OpenStudy (anonymous):

Subtract. simplify by removing a factor of 1 when possible. (10pd)/(p^2-d^2)-(p-d)/(p+d)

myininaya (myininaya):

\[\frac{10pd}{p^2-d^2}-\frac{p-d}{p+d}=\frac{10dp}{(p-q)(p+q)}-\frac{p-d}{p+d}\]

myininaya (myininaya):

find a common denominator

myininaya (myininaya):

then combine fractions

myininaya (myininaya):

multiply 2nd fraction by (p-d)/(p-d)

myininaya (myininaya):

let me know if you have anymore trouble

OpenStudy (anonymous):

wuts the answer?

myininaya (myininaya):

have you multiplied the 2nd fraction by (p-d)/(p-d) yet?

myininaya (myininaya):

now the denominators are the same so the numerator is 10dp-(p-d)(p-d)

myininaya (myininaya):

simplify the numerator i gave you and let me know what you get i will check it

OpenStudy (anonymous):

i got 12dp-p^2-d^2

myininaya (myininaya):

no thats not right (p-d)(p-d)=p^2-dp-dp+d^2=p^2-2pd+d^2

myininaya (myininaya):

so the numerator is 12dp-(p^2-2dp+d^2)

myininaya (myininaya):

now distribute the negative 1 and combine like terms and see if you can factor and cancel anything

OpenStudy (anonymous):

14dp-p^2-d^2

myininaya (myininaya):

yes very good

myininaya (myininaya):

thats the numerator

myininaya (myininaya):

and the denominator was (p-d)(p+d)

OpenStudy (anonymous):

ok so the answer is... (14dp-p^2-d^2)/(p-d)(p+d)?

myininaya (myininaya):

hey i made a type-o and i didn't realize that was a 10 not a 12 do you know what i'm talking about?

OpenStudy (anonymous):

not realy.

myininaya (myininaya):

10dp-(p^2-2dp+d^2) 10dp-p^2+2pd-d^2 12dp-p^2+d^2 is numerator ----------------------------------------- above you have in your problem 10dp not 12dp

OpenStudy (anonymous):

????????????

myininaya (myininaya):

your problem has 10dp not 12dp its the numerator of the first fraction do you still not see it?

OpenStudy (anonymous):

ok i think i see wut ur talkin about

myininaya (myininaya):

\[\frac{10dp}{p^2-d^2}-\frac{p-d}{p+d}\] this is your problem see its 10dp not 12dp so the numerator was is 10dp-(p-d)(p-d) which we already simpflied

OpenStudy (anonymous):

oooo kk

myininaya (myininaya):

and i'm saying i made a type-o and wrote 12dp-(p-d)(p-d)

myininaya (myininaya):

it is 10dp-(p-d)(p-d)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so the answer is (12dp-p^2-d^2)/(p-d)(p+d)?

OpenStudy (anonymous):

IS THAT THE RIGHT ANSWER?